OpenStudy (anonymous):

What is a factor of 4x^2 + 36x + 72

OpenStudy (anonymous):

Could someone explain step by step how to do this?

OpenStudy (ness9630):

Start off by factoring out a 4.

OpenStudy (anonymous):

This leaves x^2 + 9x + 18

OpenStudy (anonymous):

Well 4(x^2 + 9x + 18)

OpenStudy (ness9630):

Correct, now factor what's inside, and to be correct it'd be \[\LARGE 4(x^2+9x+18)=0\]

OpenStudy (anonymous):

(x + 3)(x+6)?

OpenStudy (anonymous):

But the problem is none of that looks like the answer

OpenStudy (anonymous):

Well, the given answers

OpenStudy (ness9630):

We're not finished: Now divide everything by 4 to leave: (x+3)(x+6)=0 Now solve separately: x+3=0 x+6=0

OpenStudy (anonymous):

Oh. Wow. Thanks :) So, x = - 3 and x = -6

OpenStudy (anonymous):

But there should be only one answer? So I'm guessing I did something wrong.

OpenStudy (ness9630):

Correct, and you're welcome \[\Huge\color{green}{\ddot\smile}\]

OpenStudy (ness9630):

What are your answer choices?

OpenStudy (anonymous):

One of them is x -6, so I'm guessing that is it :) It only ask for one factor.

OpenStudy (anonymous):

THANK YOU :)

OpenStudy (ness9630):

That's strange, there should be two solutions..

OpenStudy (anonymous):

Yeah, it sometimes has errors.

OpenStudy (anonymous):

COuld you help me with this one? What is a factor of x^2 - 6x -27

OpenStudy (anonymous):

Wouldn't it be (x-9)(x+3)

OpenStudy (ness9630):

Yup!

OpenStudy (anonymous):

But it doesn't list it

OpenStudy (anonymous):

x + 3 x + 9 x − 1

OpenStudy (anonymous):

And it does have None of the Above, but could it be x +3? Since that is a factor of it...?

OpenStudy (ness9630):

When factoring, you usually have TWO solutions. I know for a fact that both problems we did had two solutions. I would pick "None of The Above."

OpenStudy (ness9630):

Unless there was a mistake with the answer choices.

OpenStudy (anonymous):

It is just asking for one factor. Could that be it?

OpenStudy (ness9630):

Oh, one factor. That clears everything up.. so yes.

OpenStudy (anonymous):

Ha ha, sorry, my mistake

OpenStudy (ness9630):

You can always check your solution too, by using synthetic division.

OpenStudy (anonymous):

What about this one: 48x^2 - 26x + 3? It doesn't have any factors?

OpenStudy (anonymous):

I haven't learned synthetic division yet

OpenStudy (ness9630):

It's pretty simple, it's like upside down division.

OpenStudy (ness9630):

Here, as an example we will use the previous problem: